# Whitehead’s problem and condensed mathematics

### v3

@@ 8. Topological spaces in forcing extensions
Suppose that $(X,\tau)$ is a compact Hausdorff space and $W$ is an outer model of $V$. Then there is a canonical way to interpret $(X,\tau)$ as a compact Hausdorff space $(\hat{X},\hat{\tau})$ in $W$. First, let $\tau_{c}$ denote the collection of all closed subsets of $X$ (in $V$). In $W$, let $\hat{X}$ denote the collection of all maximal filters on $\tau_{c}\setminus\{\emptyset\}$. In other words, $\hat{X}$ is the collection of all maximal collections of elements of $\tau_{c}$ with the finite intersection property.
There is a natural map $\pi:X\rightarrow\hat{X}$ defined by letting $\pi(x):=\{D\in\tau_{c}\mid x\in D\}$ for all $x\in X$. We also define $\pi:\tau\rightarrow\mathscr{P}(\hat{X})$ by letting $\pi(U):=\{F\in\hat{X}\mid(X\setminus U)\notin F\}$. Then let $\hat{\tau}$ be the topology on $\hat{X}$ generated by $\pi[\tau]$.
@@
The space $(\hat{X},\hat{\tau})$ identified above, together with the associated maps $\pi:X\rightarrow\hat{X}$ and $\pi:\tau\rightarrow\hat{\tau}$ is precisely the interpretation of $(X,\tau)$ in $W$, in the sense of [36]. In particular, this implies the following properties, which can also readily be directly verified:
  • $(\hat{X},\hat{\tau})$ is compact in $W$;
@@
Now suppose that $\mathbb{B}$ is a complete Boolean algebra and $S=S(\mathbb{B})$ is its Stone space. We now show how to identify the space $(\hat{X},\hat{\tau})$, as computed in the forcing extension by $\mathbb{B}$, with an appropriate quotient of the space $C(S,X)^{V}$, equipped with the compact-open topology. In fact, we will prove something more general.
We first establish a correspondence between elements of $C(S,X)$ and $\mathbb{B}$-names for points in $\hat{X}$. [^3] Recent work in this direction, focusing especially on the case in which $X$ is a (compact) Polish space, can be found in [32]; many of the basic results we present here are implicit in that work. In the forward direction, fix a function $f\in C(S,X)$, and define a $\mathbb{B}$-name $\dot{F}_{f}$ for a subset of $\tau_{c}$ as follows. For every $D\in\tau_{c}$ and every $b\in\mathbb{B}$, put $(\check{D},b)$ in $\dot{F}_{f}$ if and only if $N_{b}\subseteq f^{-1}[D]$.
@@
The fact that $\dot{F}_{f}$ is forced to be a filter is immediate from the definition. To show that it is forced to be maximal, fix an arbitrary $C\in\tau_{c}$. It suffices to show that the following set is dense in $\mathbb{B}$:
$E_{C}:=\{b\in\mathbb{B}\mid N_{b}\subseteq f^{-1}[C]\text{ {or} }\exists D\in\tau_{c}\,[D\cap C=\emptyset\wedge N_{b}\subseteq f^{-1}[D]]\}.$
To this end, fix an arbitrary $b_{0}\in\mathbb{B}$, and assume that $N_{b_{0}}\not\subseteq f^{-1}[C]$. Fix $\mathcal{U}\in N_{b_{0}}$ such that $f(\mathcal{U})\in X\setminus C$. Since $X$ is compact, it is regular, so we can find an open set $O\in\tau$ such that $\mathcal{U}\in O\subseteq\mathrm{cl}(O)\subseteq X\setminus C$. By the continuity of $f$, we can find $b\leq b_{0}$ such that $\mathcal{U}\in N_{b}$ and $N_{b}\subseteq f^{-1}[O]$. Then $b\in E_{C}$, as desired. ∎
@@
Thus, $\dot{F}_{\mathcal{U}}$ has the finite intersection property; since $X$ is compact, it follows that $\bigcap\dot{F}_{\mathcal{U}}$ is nonempty. Now suppose for sake of contradiction that $x_{0}$ and $x_{1}$ are distinct elements of $\dot{F}_{\mathcal{U}}$. Since $X$ is Hausdorff, we can find disjoint open neighborhoods $U_{0}$ and $U_{1}$ of $x_{0}$ and $x_{1}$, respectively. For $i<2$, let $D_{i}:=X\setminus U_{i}$. Since $D_{0}\cup D_{1}=X$ and $\dot{F}$ is forced to be a maximal filter on $\tau_{c}\setminus\{\emptyset\}$, a routine argument shows that $\llbracket D_{0}\in\dot{F}\rrbracket_{\mathbb{B}}\vee\llbracket D_{1}\in\dot{F}\rrbracket_{\mathbb{B}}=1_{\mathbb{B}}$. Therefore, since $\mathcal{U}$ is an ultrafilter, there must be $i<2$ such that $\llbracket D_{i}\in\dot{F}\rrbracket_{\mathbb{B}}\in\mathcal{U}$, and hence $D_{i}\in\dot{F}_{\mathcal{U}}$. But then $x_{1-i}\notin D_{i}$ and $\bigcap\dot{F}_{\mathcal{U}}\subseteq D_{i}$, so $x_{1-i}\notin\bigcap\dot{F}_{\mathcal{U}}$, which is a contradiction. It follows that $\bigcap\dot{F}_{\mathcal{U}}$ contains precisely one point. ∎
@@
It is routine to verify that this is a reflexive duality: Given $g\in C(S,X)$, we have $f_{\dot{F}_{g}}=g$ and, given a $\mathbb{B}$-name $\dot{E}$ for a maximal filter on $\tau_{c}$, we have $\Vdash_{\mathbb{B}}\dot{E}=\dot{F}_{f_{\dot{E}}}$. Also, this story can clearly be relativized below a given condition $b\in\mathbb{B}$: functions in $C(N_{b},X)$ correspond to $\mathbb{B}$-names $\dot{F}$ such that $b\Vdash_{\mathbb{B}}``\dot{F}\text{ is a maximal filter on }\tau_{c}"$.
@@
First, we let the underlying set, $X_{\mathcal{K}}$, be the direct limit (i.e., colimit) of the system $\langle\hat{K},\hat{\iota}_{KK^{\prime}}\mid K,K^{\prime}\in\mathcal{K}(X),\ K\subseteq K^{\prime}\rangle$. Formally, $X_{\mathcal{K}}$ consists of equivalence classes of the form $[(x,K)]$, where $K\in\mathcal{K}(X)$ and $x\in\hat{K}$. Given an open set $U\in\tau$ and $K\in\mathcal{K}(X)$, let $\hat{U}_{K}$ denote its interpretation in $\hat{\tau}_{K}$. We then define $\hat{U}_{\mathcal{K}}\subseteq\hat{X}_{\mathcal{K}}$ as follows: given $K\in\mathcal{K}(X)$ and $x\in\hat{K}$, put $[(x,K)]\in\hat{U}_{\mathcal{K}}$ if and only if $x\in\hat{U}_{K}$. It is routine to check that this is well-defined. Now let $\hat{\tau}_{\mathcal{K}}$ be the topology generated by $\{\hat{U}_{\mathcal{K}}\mid U\in\tau\}$.
@@
It suffices to show that, for every $x\in({{}^{\omega}}\omega)^{V}$, $O\cap\hat{K}_{x}\in\hat{\tau}_{K_{x}}$. Fix such an $x$, and let $A:=\{n<\omega\mid d(n)<x(n)\}$. By choice of $d$, $A$ is finite. Then, in $V$, the set $U=\{y\in{{}^{\omega}}\omega\mid\forall n\in A~{}y(n)\neq d(n)\}$ is in $\tau$, and we have $O\cap\hat{K}_{x}=\hat{U}_{K_{x}}$. ∎
@@
Then, in $W$, we have $y\in O\setminus\hat{U}_{\mathcal{K}}$, so $\hat{U}_{\mathcal{K}}\not\subseteq O$. ∎
@@
We now show that $\hat{X}_{\mathcal{K}}$ is homeomorphic to a natural quotient of $C(S,X)^{V}$ in the extension by $\mathbb{B}$. Recall that we endow $C(S,X)^{V}$ with the compact-open topology. Now suppose that $G$ is a generic ultrafilter on $\mathbb{B}$, and, in $V[G]$, define an equivalence relation $\sim_{G}$ on $C(S,X)^{V}$ by letting $g\sim_{G}h$ if and only if there is $b\in G$ such that $g\restriction N_{b}=h\restriction N_{b}$. Let $C(S,X)^{V}/G$ denote the quotient space with respect to this equivalence relation. For each $U\in\tau$, define a set $N_{U}\subseteq C(S,X)^{V}/G$ as follows: for all $g\in C(S,X)^{V}$, put $[g]\in N_{U}$ if and only if there is $b\in G$ such that $g[N_{b}]\subseteq U$; note that this is well-defined.
@@
It is immediate from the definition that each $N_{U}$ is open. Now let $\hat{O}\subseteq C(S,X)^{V}/G$ be open, and fix $g\in C(S,X)^{V}$ such that $[g]\in\hat{O}$. Let $O=\{h\in C(S,X)^{V}\mid[h]\in\hat{O}\}$. We can then find a compact $K\subseteq S$ and an open set $U\in\tau$ such that $g[K]\subseteq U$ and $\{h\in C(S,X)^{V}\mid h[K]\subseteq U\}\subseteq O$. Since $K$ is compact and $g$ is continuous, we can find a finite collection $b_{0},\ldots,b_{n-1}$ from $\mathbb{B}$ such that $K\subseteq\bigcup\{N_{b_{i}}\mid i<n\}$ and, for all $i<n$, we have $g[N_{b_{i}}]\subseteq U$. Letting $b:=\bigvee\{b_{i}\mid i<n\}$, we get $K\subseteq N_{b}$ and $g[N_{b}]\subseteq U$.
Suppose first that $b\in G$. In this case, we claim that $N_{U}\subseteq\hat{O}$. To see this, fix $h\in C(S,X)^{V}$ such that $[h]\in N_{U}$. We can therefore find $c\in G$ such that $h[N_{c}]\subseteq U$. Define $h^{\prime}\in C(S,X)^{V}$ by letting $h^{\prime}\restriction N_{c}=h\restriction N_{c}$ and letting $h^{\prime}\restriction(S\setminus N_{c})$ be constant, taking an arbitrary value in $U$. Then $h^{\prime}[K]\subseteq U$, so $[h^{\prime}]\in\hat{O}$. Moreover, since $c\in G$, we have $h^{\prime}\sim_{G}h$, so $[h^{\prime}]=[h]$. Thus, $N_{U}\subseteq\hat{O}$.
Suppose next that $b\notin G$. In this case, we claim that $\hat{O}$ is actually all of $C(S,X)^{V}/G$. To see this, let $h\in C(S,X)^{V}$ be arbitrary. Define $h^{\prime}\in C(S,X)^{V}$ by letting $h^{\prime}\restriction N_{b}=g\restriction N_{b}$ and $h^{\prime}\restriction(S\setminus N_{b})=h\restriction(S\setminus N_{b})$. Then $h^{\prime}[N_{b}]=g[N_{b}]\subseteq U$, so $[h^{\prime}]\in\hat{O}$. Since $b\notin G$, we have $[h^{\prime}]=[h]$, so $[h]\in\hat{O}$. ∎
@@
$\{b\in\mathbb{B}\mid g\restriction N_{b}=h\restriction N_{b}\text{ {or} }g[N_{b}]\cap h[N_{b}]=\emptyset\}$
is dense in $\mathbb{B}$. The genericity of $G$ then yields the injectivity of $k$. To show that $k$ is surjective, fix an element $[(F,K)]\in\hat{X}_{\mathcal{K}}$, and let $\dot{F}$ be a $\mathbb{B}$-name for $F$ that is forced by $1_{\mathbb{B}}$ to be a maximal filter on $(\tau_{K})_{c}$. Then $f_{\dot{F}}$, as defined at the beginning of this section, is in $C(S,X)^{V}$, and it is routine to verify that $k([f_{\dot{F}}])=[(F,K)]$. ∎
@@
Suppose that $g\in C(S,X)^{V}$ and $[g]\in N_{U}$. Let $K:=g[S]$, and fix $b\in G$ such that $g[N_{b}]\subseteq U$. Then, by construction, we have $(K\setminus U)\notin F_{g}$, so $k([g])=[(F_{g},K)]\in\hat{U}_{\mathcal{K}}$. Conversely, suppose that $g\in C(S,X)^{V}$ and $[g]\notin N_{U}$. Again, let $K:=g[S]$. By familiar arguments, the following set is dense in $\mathbb{B}$ (in $V$):
$\{b\in\mathbb{B}\mid g[N_{b}]\subseteq U\text{ {or} }g[N_{b}]\subseteq K\setminus U\}.$
Since $[g]\notin N_{U}$, the genericity of $G$ implies that $K\setminus U\in F_{g}$, so $k([g])=[(F_{g},K))]\notin\hat{U}_{\mathcal{K}}$. ∎
@@
Second, write $cg$ for the left adjoint of the inclusion of the category of compactly generated topological spaces into the category of topological spaces; recall also that the embedding $X\mapsto\underline{X}$ of the category of $T_{1}$ topological spaces into Cond possesses a left adjoint whose counit is given by $X\mapsto cg(X)$ (modulo a small technical point; see [26, pp. 8-9 and Appendix to Lecture II]). Observe next that in the $\mathbb{B}=\{0,1\}$ (and hence $S(\mathbb{B})=*$) instance of Theorem 8.11, $\hat{X}_{\mathcal{K}}$ is homeomorphic to $X$. Had we endowed it instead with the $\hat{\tau}^{\lim}_{\mathcal{K}}$ topology considered above, it would be homeomorphic to $cg(X)$, and the $\hat{\tau}^{\lim}_{\mathcal{K}}$ topology corresponds more generally to that given by post-composing the conversion recorded in Theorem 8.11 with the functor $cg$.
Third, we note that the operation $X\mapsto\hat{X}_{\mathcal{K}}$ commutes with products. Namely, if $X_{i}$ is a Hausdorff space for every $i\in I$, then $(\widehat{\prod_{i\in I}X_{i}})_{\mathcal{K}}\cong\prod_{i\in I}(\hat{X}_{i})_{\mathcal{K}}$. This is because the collection
@@
Fourth, even though Theorem 8.11 holds for all Hausdorff spaces $X$, there is an important sense in which the correspondence is meaningfully stronger in case $X$ is compact. In particular, let $\dot{k}$ be the $\mathbb{B}$-name for the homeomorphism we constructed to witness Theorem 8.11. If $X$ is compact, then, for every $\mathbb{B}$-name $\dot{x}$ for an element of $\hat{X}_{\mathcal{K}}(=\hat{X})$, there is a $g\in C(S,X)$ such that $1_{\mathbb{B}}\Vdash\dot{k}([g])=\dot{x}$. In particular, in this case we can take $\dot{x}$ to be a $\mathbb{B}$-name for a maximal filter on $\tau_{c}$, then then let $g=f_{\dot{x}}$, as defined at the outset of this section. However, in general this may not be the case, and there may be $\mathbb{B}$-names $\dot{x}$ for elements of $\hat{X}_{\mathcal{K}}$ for which there is no $g\in C(S,X)$ such that $1_{\mathbb{B}}\Vdash\dot{k}([g])=\dot{x}$. For such $\dot{x}$, one may need to work below some condition $b\in\mathbb{B}$ that forces $\dot{x}$ to lie inside the interpretation of some particular $K\in\mathcal{K}(X)$; once one does so, one may then find $g\in C(S,X)$ such that $b\Vdash\dot{k}([g])=\dot{x}$. This issue will lead to some slight complications and asymmetries in the next section.

### v5

@@ 8. Topological spaces in forcing extensions
[…]
@@
[…]
@@
[…]
We first establish a correspondence between elements of $C(S,X)$ and $\mathbb{B}$-names for points in $\hat{X}$. [^4] Recent work in this direction, focusing especially on the case in which $X$ is a (compact) Polish space, can be found in [34]; many of the basic results we present here are implicit in that work. In the forward direction, fix a function $f\in C(S,X)$, and define a $\mathbb{B}$-name $\dot{F}_{f}$ for a subset of $\tau_{c}$ as follows. For every $D\in\tau_{c}$ and every $b\in\mathbb{B}^{+}$, put $(\check{D},b)$ in $\dot{F}_{f}$ if and only if $N_{b}\subseteq f^{-1}[D]$.
@@
The fact that $\dot{F}_{f}$ is forced to be a filter is immediate from the definition. To show that it is forced to be maximal, fix an arbitrary $C\in\tau_{c}$. It suffices to show that the following set is dense in $\mathbb{B}^{+}$:
$E_{C}:=\{b\in\mathbb{B}^{+}\mid N_{b}\subseteq f^{-1}[C]\text{ {or} }\exists D\in\tau_{c}\,[D\cap C=\emptyset\wedge N_{b}\subseteq f^{-1}[D]]\}.$
To this end, fix an arbitrary $b_{0}\in\mathbb{B}^{+}$, and assume that $N_{b_{0}}\not\subseteq f^{-1}[C]$. Fix $\mathcal{U}\in N_{b_{0}}$ such that $f(\mathcal{U})\in X\setminus C$. Since $X$ is compact, it is regular, so we can find an open set $O\in\tau$ such that $\mathcal{U}\in O\subseteq\mathrm{cl}(O)\subseteq X\setminus C$. By the continuity of $f$, we can find $b\leq b_{0}$ such that $\mathcal{U}\in N_{b}$ and $N_{b}\subseteq f^{-1}[O]$. Then $b\in E_{C}$, as desired. ∎
@@
Thus, $\dot{F}_{\mathcal{U}}$ has the finite intersection property; since $X$ is compact, it follows that $\bigcap\dot{F}_{\mathcal{U}}$ is nonempty. Now suppose for the sake of contradiction that $x_{0}$ and $x_{1}$ are distinct elements of $\bigcap\dot{F}_{\mathcal{U}}$. Since $X$ is Hausdorff, we can find disjoint open neighborhoods $U_{0}$ and $U_{1}$ of $x_{0}$ and $x_{1}$, respectively. For $i<2$, let $D_{i}:=X\setminus U_{i}$. Since $D_{0}\cup D_{1}=X$ and $\dot{F}$ is forced to be a maximal filter on $\tau_{c}\setminus\{\emptyset\}$, a routine argument shows that $\llbracket D_{0}\in\dot{F}\rrbracket_{\mathbb{B}}\vee\llbracket D_{1}\in\dot{F}\rrbracket_{\mathbb{B}}=1_{\mathbb{B}}$. Therefore, since $\mathcal{U}$ is an ultrafilter, there must be $i<2$ such that $\llbracket D_{i}\in\dot{F}\rrbracket_{\mathbb{B}}\in\mathcal{U}$, and hence $D_{i}\in\dot{F}_{\mathcal{U}}$. But then $x_{i}\notin D_{i}$ and $\bigcap\dot{F}_{\mathcal{U}}\subseteq D_{i}$, so $x_{i}\notin\bigcap\dot{F}_{\mathcal{U}}$, which is a contradiction. It follows that $\bigcap\dot{F}_{\mathcal{U}}$ contains precisely one point. ∎
@@
It is routine to verify that this is a reflexive duality: Given $g\in C(S,X)$, we have $f_{\dot{F}_{g}}=g$ and, given a $\mathbb{B}$-name $\dot{E}$ for a maximal filter on $\tau_{c}$, we have $\Vdash_{\mathbb{B}}\dot{E}=\dot{F}_{f_{\dot{E}}}$. Also, this story can clearly be relativized below a given condition $b\in\mathbb{B}^{+}$: functions in $C(N_{b},X)$ correspond to $\mathbb{B}$-names $\dot{F}$ such that $b\Vdash_{\mathbb{B}}``\dot{F}\text{ is a maximal filter on }\tau_{c}$".
@@
First, we let the underlying set, $X_{\mathcal{K}}$, be the direct limit (i.e., colimit) of the system $\langle\hat{K},\hat{\iota}_{KK^{\prime}}\mid K,K^{\prime}\in\mathcal{K}(X),\ K\subseteq K^{\prime}\rangle$. Formally, $X_{\mathcal{K}}$ consists of equivalence classes of the form $[(x,K)]$, where $K\in\mathcal{K}(X)$ and $x\in\hat{K}$. Given an open set $U\in\tau$ and $K\in\mathcal{K}(X)$, let $\hat{U}_{K}$ denote its interpretation in $\hat{\tau}_{K}$. We then define $\hat{U}_{\mathcal{K}}\subseteq\hat{X}_{\mathcal{K}}$ as follows: given $K\in\mathcal{K}(X)$ and $x\in\hat{K}$, put $(x,K)]\in\hat{U}_{\mathcal{K}}$ if and only if $x\in\hat{U}_{K}$. It is routine to check that this is well-defined. Now let $\hat{\tau}_{\mathcal{K}}$ be the topology generated by $\{\hat{U}_{\mathcal{K}}\mid U\in\tau\}$.
@@
It suffices to show that, for every $x\in({{}^{\omega}}\omega)^{V}$, $O\cap\hat{K}_{x}\in\hat{\tau}_{K_{x}}$. Fix such an $x$, and let $A:=\{n<\omega\mid d(n)<x(n)\}$. By the choice of $d$, AA is finite. Then, in $V$, the set $U=\{y\in{{}^{\omega}}\omega\mid\forall n\in A~y(n)\neq d(n)\}$ is in $\tau$, and we have $O\cap\hat{K}_{x}=\hat{U}_{K_{x}}$. ∎
@@
Then, in $W$, we have $y\in\hat{U}_{\mathcal{K}}\setminus O$, so $\hat{U}_{\mathcal{K}}\not\subseteq O$. ∎
@@
We now show that $\hat{X}_{\mathcal{K}}$ is homeomorphic to a natural quotient of $C(S,X)^{V}$ in the extension by $\mathbb{B}$. Recall that we endow $C(S,X)^{V}$ with the compact-open topology. Now suppose that $G$ is a generic ultrafilter on $\mathbb{B}$, and, in $V[G]$, define an equivalence relation $\sim_{G}$ on $C(S,X)^{V}$ by letting $g\sim_{G}h$ if and only if there is $b\in G$ such that $g\restriction N_{b}=h\restriction N_{b}$. Let $C(S,X)^{V}/G$ denote the quotient space with respect to this equivalence relation. For each $U\in\tau$, define a set $N_{U}\subseteq C(S,X)^{V}/G$ as follows: for all $g\in C(S,X)^{V}$, put $g]\in N_{U}$ if and only if there is $b\in G$ such that $g[N_{b}]\subseteq U$; note that this is well-defined.
@@
It is immediate from the definition that each $N_{U}$ is open. Now let $\hat{O}\subseteq C(S,X)^{V}/G$ be open, and fix $g\in C(S,X)^{V}$ such that $g]\in\hat{O}$. Let $O=\{h\in C(S,X)^{V}\mid[h]\in\hat{O}\}$. We can then find a compact $K\subseteq S$ and an open set $U\in\tau$ such that $g[K]\subseteq U$ and $\{h\in C(S,X)^{V}\mid h[K]\subseteq U\}\subseteq O$. Since $K$ is compact and $g$ is continuous, we can find a finite collection $b_{0},\ldots,b_{n-1}$ from $\mathbb{B}^{+}$ such that $K\subseteq\bigcup\{N_{b_{i}}\mid i<n\}$ and, for all $i<n$, we have $g[N_{b_{i}}]\subseteq U$. Letting $b:=\bigvee\{b_{i}\mid i<n\}$, we get $K\subseteq N_{b}$ and $g[N_{b}]\subseteq U$.
Suppose first that $b\in G$. In this case, we claim that $N_{U}\subseteq\hat{O}$. To see this, fix $h\in C(S,X)^{V}$ such that $h]\in N_{U}$. We can therefore find $c\in G$ such that $h[N_{c}]\subseteq U$. Define $h^{\prime}\in C(S,X)^{V}$ by letting $h^{\prime}\restriction N_{c}=h\restriction N_{c}$ and letting $h^{\prime}\restriction(S\setminus N_{c})$ be constant, taking an arbitrary value in $U$. Then $h^{\prime}[K]\subseteq U$, so $h^{\prime}]\in\hat{O}$. Moreover, since $c\in G$, we have $h^{\prime}\sim_{G}h$, so $h^{\prime}]=[h]$. Thus, $N_{U}\subseteq\hat{O}$.
Suppose next that $b\notin G$. In this case, we claim that $\hat{O}$ is actually all of $C(S,X)^{V}/G$. To see this, let $h\in C(S,X)^{V}$ be arbitrary. Define $h^{\prime}\in C(S,X)^{V}$ by letting $h^{\prime}\restriction N_{b}=g\restriction N_{b}$ and $h^{\prime}\restriction(S\setminus N_{b})=h\restriction(S\setminus N_{b})$. Then $h^{\prime}[N_{b}]=g[N_{b}]\subseteq U$, so $h^{\prime}]\in\hat{O}$. Since $b\notin G$, we have $h^{\prime}]=[h]$, so $h]\in\hat{O}$. ∎
@@
$\{b\in\mathbb{B}^{+}\mid g\restriction N_{b}=h\restriction N_{b}\text{ {or} }g[N_{b}]\cap h[N_{b}]=\emptyset\}$
is dense in $\mathbb{B}^{+}$. The genericity of $G$ then yields the injectivity of $k$. To show that $k$ is surjective, fix an element $(F,K)]\in\hat{X}_{\mathcal{K}}$, and let $\dot{F}$ be a $\mathbb{B}$-name for $F$ that is forced by $1_{\mathbb{B}}$ to be a maximal filter on $(\tau_{K})_{c}$. Then $f_{\dot{F}}$, as defined at the beginning of this section, is in $C(S,X)^{V}$, and it is routine to verify that $k([f_{\dot{F}}])=[(F,K)]$. ∎
@@
Suppose that $g\in C(S,X)^{V}$ and $g]\in N_{U}$. Let $K:=g[S]$, and fix $b\in G$ such that $g[N_{b}]\subseteq U$. Then, by construction, we have $(K\setminus U)\notin F_{g}$, so $k([g])=[(F_{g},K)]\in\hat{U}_{\mathcal{K}}$. Conversely, suppose that $g\in C(S,X)^{V}$ and $g]\notin N_{U}$. Again, let $K:=g[S]$. By familiar arguments, the following set is dense in $\mathbb{B}^{+}$ (in $V$):
$\{b\in\mathbb{B}^{+}\mid g[N_{b}]\subseteq U\text{ {or} }g[N_{b}]\subseteq K\setminus U\}.$
Since $g]\notin N_{U}$, the genericity of $G$ implies that $K\setminus U\in F_{g}$, so $k([g])=[(F_{g},K))]\notin\hat{U}_{\mathcal{K}}$. ∎
@@
Second, write $cg$ for the left adjoint of the inclusion of the category of compactly generated topological spaces into the category of topological spaces; recall also that the embedding $X\mapsto\underline{X}$ of the category of $T_{1}$ topological spaces into Cond possesses a left adjoint whose counit is given by $X\mapsto cg(X)$ (modulo a small technical point; see [28, pp. 8-9 and Appendix to Lecture II]). Observe next that in the $\mathbb{B}=\{0,1\}$ (and hence $S(\mathbb{B})=*$) instance of Theorem 8.11, $\hat{X}_{\mathcal{K}}$ is homeomorphic to $X$. Had we endowed it instead with the $\hat{\tau}^{\lim}_{\mathcal{K}}$ topology considered above, it would be homeomorphic to $cg(X)$, and the $\hat{\tau}^{\lim}_{\mathcal{K}}$ topology corresponds more generally to that given by post-composing the conversion recorded in Theorem 8.11 with the functor $cg$.
Third, we note that the operation $X\mapsto\hat{X}_{\mathcal{K}}$ commutes with products. Namely, if $I$ is an index set and $X_{i}$ is a Hausdorff space for every $i\in I$, then $(\widehat{\prod_{i\in I}X_{i}})_{\mathcal{K}}\cong\prod_{i\in I}(\hat{X}_{i})_{\mathcal{K}}$. This is because the collection
@@
[…]